# Holographic analysis of near-conformal dynamics and light dilaton

### v1

@@ Appendix B Invariant fluctuations and fluctuation equations
Here $i,j$ are the flavor indices. We only consider flavor independent case so we can take the flavor background and fluctuations to be proportional to the unit matrix. We focus here on the scalar (spin zero) sector. This sector has two dynamical fields. It is convenient to use the diffeomorphism invariant combinations. To do this we define
@@
The first order fluctuation action around the background solution is always a boundary term. It may be written as
@@
The renormalized second order action is written as
@@ Appendix D Expansions near the UV boundary
In order to analyze the UV divergence of the action, we need the expansions of the background and the fluctuations near the boundary. We start from the background. By using the equations of motion from Appendix C, it is found to obey the following asymptotics at the boundary:
@@
In the standard gauge (i.e., $\hat{\phi}=0$ and $\partial^{\mu}\hat{A}_{\mu}=0$), we find for the gauge noninvariant functions
@@
The functions $F_{1,2}$ were normalized such that $F_{1}(0)=1=F_{2}(0)$. We omitted corrections suppressed by integer powers of $z$ to each term which also contain nontrivial dependence on space-time coordinates. Notice that there are three independent sources ($\chi_{-}$, $\psi_{-}$, $\tau_{-}$) but only one vev coefficient $\xi_{+}$.
@@ Appendix E Holographic renormalization
After adding the counterterm contributions to (166), the first order fluctuation term becomes
@@
After adding the second order contribution in fluctuations $S^{(2)}_{\mathrm{ct}}$ from the counterterm (196), the second order action (167) simplifies to
@@ E.1 Correlators
where the unnormalized correlator $\widetilde{G}(q)$ likewise only depends on the momentum through $q^{2}$. Fourier transforming this back to coordinate space gives
@@
where we reinstated the dilaton contribution. Notice that it looks like setting here $\Delta_{g}=4$ would just eliminate the gluon contribution. This is however not that straightforward because this value is a special case. In particular, a constant source solution of $\phi$ will not represent a good leading RG flow end at $\phi=0$. Higher order corrections of the dilaton potential will modify the flow of source as we will se below in Appendix F.
@@
$\langle\bar{\psi}\psi\rangle$ $=$ $i\frac{\delta}{\delta\chi_{-}(x)}\log Z_{\mathrm{QCD}}[\chi_{-}]\Big{|}_{\chi_{-}=0}=i\frac{\delta}{\delta\chi_{-}(x)}\exp(iS^{(1)}_{\mathrm{ren}})\Big{|}_{\chi_{-}=0}$ (213)
@@ Appendix F Holographic RG flow and the ward identity
In this appendix sketch how to obtain an expression of the correlator $\langle\theta^{\mu}_{\mu}(x)\rangle$ in terms of the beta function and the gluon condensate. Following Gursoy:2007cb , we identify the ’t Hooft coupling as $\lambda=e^{\phi}$ and the logarithmic of the energy scale as $A(r)\approx-\log r$. Therefore
@@
$\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{4\ell^{3}}{\kappa^{2}N_{c}}\left(\Delta_{g}-2\right)\phi_{v}\epsilon^{\Delta_{g}-4}e^{\phi_{\mathrm{UV}}}$ (215)
@@
$-\frac{2\ell^{3}}{\kappa^{2}}(\Delta_{g}-2)(4-\Delta_{g})\phi_{s}\phi_{v}\approx\frac{N_{c}}{2\lambda^{2}}\beta(\epsilon)\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle$ (216)
@@
$\langle\theta^{\mu}_{\mu}\rangle=(4-\Delta)m_{q}\langle\bar{\psi}\psi\rangle+\frac{N_{c}}{2\lambda^{2}}\beta(\epsilon)\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\ .$ (217)
@@ F.1 Different dilation potential examples
After considering the boundary conditions from appendix D.1 and inserting the Ansatz for the fields the first order action becomes
@@
$\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{8\ell^{3}\phi_{v}}{\kappa^{2}N_{c}}\left[\left(-\log(\epsilon)\right)^{\frac{n-1}{n-2}}-\frac{n^{2}\ell^{2}}{32}\lambda_{g}\phi^{n-1}_{s}\phi_{v}\right]e^{\phi_{\mathrm{UV}}}$ (227)
@@
$\frac{N_{c}}{4\lambda^{2}}\beta(\epsilon)\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{2\ell^{3}\phi_{s}\phi_{v}}{\kappa^{2}(n-2)}$ (228)
@@
After consider the boundary conditions from appendix D.1 and inserting the Ansatz for the fields the first order action becomes
@@
$\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{8\ell^{3}\phi_{v}}{\kappa^{2}N_{c}}\left[\log(r)+\phi_{v2}-\kappa_{s}\right]e^{\phi(r)}$ (237)
@@
$\frac{N_{c}}{4\lambda^{2}}\beta(\epsilon)\langle{\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}\rangle\approx-\frac{4\ell^{3}\phi_{v}}{\sqrt{3}\kappa^{2}}$ (238)
@@ Appendix G On the generality of the IR boundary conditions
In the main text, we used generic IR boundary conditions (see Eqs. (33) and (126)). The way we define the boundary conditions, in particular in the model computation of Sec. 3, may appear dubious because the definition (33) on the one hand assumes linearity of the equation of motion of $X$ but on the other hand in the IR regime $r\sim r_{\mathrm{IR}}$ the field $X$ will grow large so that nonlinearities and backreaction to the metric become important. In the analytic model per se there is no problem, as nonlinearities and backreaction are absent by definition, but we argue that the model also represents setups where these features are included. Here we show why this is the case.
We begin by considering a generic action for $X$ which we do not specify explicitly, but we assume that it is backreacted to the gravity sector and agrees with (73) when $X\ll 1$. We denote the backreacted regular solution of the potentially nonlinear action by $X_{\mathrm{nl}}$. For concrete examples (such as the model in Jarvinen:2011qe ) such a solution can be found numerically. Without picking an explicit model we cannot specify what is meant by regularity, but for the model to be consistent and able to describe walking geometries, there needs to be a way to pick a single solution such that $X$ becomes small when evolved towards the boundary, while the geometry approaches the IR fixed point in the walking regime, $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$. That is, the background takes the form of (79) while the flavor field obeys (89):
@@
We then consider a general solution of $X_{\mathrm{lin}}$ to the linearized action and in the absence of backreaction. This solution obeys [^1] Note that in the absence of backreaction, the background stay at the IR fixed point for all $r\gg r_{\mathrm{UV}}$.
@@
where we can take the UV and IR scales to be the same as for the nonlinear solution. Now we can set the IR regularity condition for the linear solution to be
@@
This condition does not determine the normalization, $X_{\mathrm{IR}}$, but in the absence of backreaction and nonlinearities this parameter scales out from the action so it affects the analysis only trivially. Note however that it is a number which is independent of the UV scale $r_{\mathrm{UV}}$, and is also determined by the definition of $r_{\mathrm{IR}}$.
One can follow similar steps to show that general IR regularity conditions for the fluctuation $\xi$ can be captured by our conditions (13). The fluctuation equations are always linear, but the general regular solution should take into account the coupling between the wave functions $\xi$ and $\zeta$ in (108) and (109) which becomes sizable for $r\sim r_{\mathrm{IR}}$. This coupling can be treated in the same way as the nonlinearities for the background solution. There is however an important complication: the coefficients (for example the counterpart of $X_{\mathrm{IR}}^{\mathrm{nl}})$ will depend on the mass or frequency of the fluctuations $\omega$. However if the mass is small, $\omega\ll 1/r_{\mathrm{IR}}$, it can be neglected in the analysis. That is, the coefficients $\mathcal{A}$, $\mathcal{B}$ in (13) can be interpreted to be fixed numbers only for light modes.

### v2

@@ Appendix B Invariant fluctuations and fluctuation equations
Here $i,j$ are the flavor indices. We only consider the flavor independent case so we can take the flavor background and fluctuations to be proportional to the unit matrix. We focus here on the scalar (spin zero) sector. This sector has two dynamical fields. It is convenient to use the diffeomorphism invariant combinations. To do this we define
@@
The first-order fluctuation action around the background solution is always a boundary term. It may be written as
@@
The renormalized second-order action is written as
@@ Appendix D Expansions near the UV boundary
In order to analyze the UV divergence of the action, we need the expansions of the background and the fluctuations near the boundary. We start from the background. By using the equations of motion from Appendix C, it is found that they obey the following asymptotics at the boundary:
@@
In the standard gauge (i.e., $\hat{\phi}=0$ and $\partial^{\mu}\hat{A}_{\mu}=0$), we find for the gauge non-invariant functions
@@
The functions $F_{1,2}$ were normalized such that $F_{1}(0)=1=F_{2}(0)$. We omitted corrections suppressed by integer powers of $z$ to each term which also contain non-trivial dependence on space-time coordinates. Notice that there are three independent sources ($\chi_{-}$, $\psi_{-}$, $\tau_{-}$) but only one vev coefficient $\xi_{+}$.
@@ Appendix E Holographic renormalization
After adding the counterterm contributions to (168), the first-order fluctuation term becomes
@@
After adding the second-order contribution in fluctuations $S^{(2)}_{\mathrm{ct}}$ from the counterterm (198), the second-order action (169) simplifies to
@@ E.1 Correlators
where the un-normalized correlator $\widetilde{G}(q)$ likewise only depends on the momentum through $q^{2}$. Fourier transforming this back to coordinate space gives
@@
where we reinstated the dilaton contribution. Notice that it looks like setting here $\Delta_{g}=4$ would just eliminate the gluon contribution. This is, however, not that straightforward because this value is a special case. In particular, a constant source solution of $\phi$ will not represent a good leading RG flow end at $\phi=0$. Higher order corrections of the dilaton potential will modify the flow of source as we will se below in Appendix F.
@@
$\langle\bar{q}q\rangle$ $=$ $i\frac{\delta}{\delta\chi_{-}(x)}\log Z_{\mathrm{QCD}}[\chi_{-}]\Big{|}_{\chi_{-}=0}=i\frac{\delta}{\delta\chi_{-}(x)}\exp(iS^{(1)}_{\mathrm{ren}})\Big{|}_{\chi_{-}=0}$ (215)
@@ Appendix F Holographic RG flow and the ward identity
In this appendix, we sketch how to obtain an expression of the correlator $\langle\theta^{\mu}_{\mu}(x)\rangle$ in terms of the beta function and the gluon condensate. Following Gursoy:2007cb , we identify the ’t Hooft coupling as $\lambda=e^{\phi}$ and the logarithmic of the energy scale as $A(r)\approx-\log r$. Therefore
@@
$\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{4\ell^{3}}{\kappa^{2}N_{c}}\left(\Delta_{g}-2\right)\phi_{v}\epsilon^{\Delta_{g}-4}e^{\phi_{\mathrm{UV}}}$ (217)
@@
$-\frac{2\ell^{3}}{\kappa^{2}}(\Delta_{g}-2)(4-\Delta_{g})\phi_{s}\phi_{v}\approx\frac{N_{c}}{2\lambda^{2}}\beta(\epsilon)\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle$ (218)
@@
$\langle\theta^{\mu}_{\mu}\rangle=(4-\Delta)m_{q}\langle\bar{q}q\rangle+\frac{N_{c}}{2\lambda^{2}}\beta(\epsilon)\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\ .$ (219)
@@ F.1 Different dilation potential examples
After considering the boundary conditions from appendix D.1 and inserting the ansatze for the fields the first-order action becomes
@@
$\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{8\ell^{3}\phi_{v}}{\kappa^{2}N_{c}}\left[\left(-\log(\epsilon)\right)^{\frac{n-1}{n-2}}-\frac{n^{2}\ell^{2}}{32}\lambda_{g}\phi^{n-1}_{s}\phi_{v}\right]e^{\phi_{\mathrm{UV}}}$ (229)
@@
$\frac{N_{c}}{4\lambda^{2}}\beta(\epsilon)\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{2\ell^{3}\phi_{s}\phi_{v}}{\kappa^{2}(n-2)}$ (230)
@@
After consider the boundary conditions from appendix D.1 and inserting the Ansatz for the fields the first-order action becomes
@@
$\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\approx\frac{8\ell^{3}\phi_{v}}{\kappa^{2}N_{c}}\left[\log(r)+\phi_{v2}-\kappa_{s}\right]e^{\phi(r)}$ (239)
@@
$\frac{N_{c}}{4\lambda^{2}}\beta(\epsilon)\langle{\mathrm{Tr}}G_{\mu\nu}G^{\mu\nu}\rangle\approx-\frac{4\ell^{3}\phi_{v}}{\sqrt{3}\kappa^{2}}$ (240)
@@ Appendix G On the generality of the IR boundary conditions
In the main text, we used generic IR boundary conditions (see Eqs. (35) and (128)). The way we define the boundary conditions, in particular in the model computation of Sec. 3, may appear dubious because the definition (35) on the one hand assumes linearity of the equation of motion of $X$ but on the other hand in the IR regime $r\sim r_{\mathrm{IR}}$ the field $X$ will grow large so that non-linearities and backreaction to the metric become important. In the analytic model per se there is no problem, as non-linearities and backreaction are absent by definition, but we argue that the model also represents setups where these features are included. Here we show why this is the case.
We begin by considering a generic action for $X$ which we do not specify explicitly, but we assume that it is backreacted to the gravity sector and agrees with (75) when $X\ll 1$. We denote the backreacted regular solution of the potentially non-linear action by $X_{\mathrm{nl}}$. For concrete examples (such as the model in Jarvinen:2011qe ) such a solution can be found numerically. Without picking an explicit model we cannot specify what is meant by regularity, but for the model to be consistent and able to describe walking geometries, there needs to be a way to pick a single solution such that $X$ becomes small when evolved towards the boundary, while the geometry approaches the IR fixed point in the walking regime, $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$. That is, the background takes the form of (81) while the flavor field obeys (91):
@@
We then consider a general solution of $X_{\mathrm{lin}}$ to the linearized action and in the absence of backreaction. This solution obeys [^1] Note that in the absence of backreaction, the background stays at the IR fixed point for all $r\gg r_{\mathrm{UV}}$.
@@
where we can take the UV and IR scales to be the same as for the non-linear solution. Now we can set the IR regularity condition for the linear solution to be
@@
This condition does not determine the normalization, $X_{\mathrm{IR}}$, but in the absence of backreaction and non-linearities this parameter scales out from the action so it affects the analysis only trivially. Note, however, that it is a number which is independent of the UV scale $r_{\mathrm{UV}}$, and is also determined by the definition of $r_{\mathrm{IR}}$.
One can follow similar steps to show that general IR regularity conditions for the fluctuation $\xi$ can be captured by our conditions (13). The fluctuation equations are always linear, but the general regular solution should take into account the coupling between the wave functions $\xi$ and $\zeta$ in (110) and (111) which becomes sizable for $r\sim r_{\mathrm{IR}}$. This coupling can be treated in the same way as the non-linearities for the background solution. There is however an important complication: the coefficients (for example the counterpart of $X_{\mathrm{IR}}^{\mathrm{nl}})$ will depend on the mass or frequency of the fluctuations $\omega$. However if the mass is small, $\omega\ll 1/r_{\mathrm{IR}}$, it can be neglected in the analysis. That is, the coefficients $\mathcal{A}$, $\mathcal{B}$ in (13) can be interpreted to be fixed numbers only for light modes.